What is the oxidation state of each element in K,Cr,0,? What is the oxidation state of each element in SO-?
Answer #1

1)

Oxidation number of K = +1

Oxidation number of O = -2

lets the oxidation number of Cr be x

use:

2* oxidation number (K) + 7* oxidation number (O) + 2* oxidation number (Cr) = net charge

2*(+1)+7*(-2)+2* x = 0

-12 + 2 * x = 0

x = 6

So oxidation number of Cr = +6

Oxidation number of each element in K2Cr2O7 is

K=+1

O=-2

Cr=+6

2)

Oxidation number of O = -2

lets the oxidation number of S be x

use:

3* oxidation number (O) + 1* oxidation number (S) = net charge

3*(-2)+1* x = -2

-6 + 1 * x = -2

x = 4

So oxidation number of S = +4

Oxidation number of each element in SO3-2 is

O=-2

S=+4